[figure: 9.7 cm]

MAGNETIC PROPERTIES AND ELECRONIC STRUCTURE OF

THE Gd(Co_{1−x}Si_{x})_{2} AND Gd(Co_{1−x}Ge_{x})_{2} LAVES PHASE COMPOUNDS

I. Debiec, G. Chelkowska Institute of Physics, University of Silesia, ul. Uniwersytecka 4, 40-007 Katowice, Poland

Results of measurements of the magnetic susceptibilty, the electrical resistivity and the crystal structure of the polycrystalline intermetallic compounds Gd(Co_{1−x}Si_{x})_{2} and Gd(Co_{1−x}Ge_{x})_{2} with x=0; 0.05; 0.1; 0.15; 0.20 are presented. The effect of the partial substitution Co by Si and Ge is similar and reflects in the change of the Curie temperatures T_{C} and in the temperature dependence of the electrical resistivity. In both cases (Si, Ge) we cannot detect the change in the lattice parameters. In the condition of the fixed volume the hybridization between the Co d states and Si(Ge) p states seems to be responsible for the observed effects, what was confirmed by XPS measurements in the valence band regions.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : G. Chelkowska

Address for correspondence : ul. Uniwersytecka 4 40-007 Katowice

Email address : gchelkow@us.edu.pl